Helen has 1/5 of her portfolio in US stocks, 1/8 of her portfolio in European stocks, and 1/10 of her portfolio in Japanese stocks. The remainder is invested in municipal bonds. What percent is invested in municipal bonds?
Answer by rfer(16322)

40/40-(1/8+1/5+1/10)=
40/40-(5/40+8/40+4/40)=
40/40-17/40=23/40 or 57.5% in municipal bonds